OpenStudy (nottim):

A boat is travelling at 18 knots on a heading of 234 degrees. The current is 8 knots, flowing from a bearing of 124 degrees. Determine the resultant velocity of the boat.

OpenStudy (nottim):

I got the speed, its 22 knots. But I didn't get the angle.

OpenStudy (nottim):

The correct angle is a bearing of 254. I got 19.9 degrees.

OpenStudy (nottim):

to get such an angle, i used sin^-1(8sin110/22.056)=Angle

OpenStudy (nottim):

to get my speed, c2=8^2-182-2(8)(18)cos110

OpenStudy (nottim):

um...anyone got a response?

OpenStudy (dumbcow):

everything looks correct, again the 19.9 refers to the angle inside the triangle the new bearing will be 234 + 19.9 = 253.9 which is slighly different from answer given ? not sure why

OpenStudy (nottim):

this txtbook rounds the answers. How do I know whether to add or subtract from the current bearing on which the angle/point lies?

OpenStudy (dumbcow):

 use the diagram